Construct a Pareto chart for the number of registered taxicabs in the selected cities.
(City)=new York, Washington,D.C, Chicago, Philadelphia, baltimore.
(Number)=11,787 8348 , 5300 , 1480 , 115
1
Expert's answer
2020-03-23T15:51:44-0400

Arrange the data from the largest to smallest according to frequency.


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 City & Number \\ \hline
 D.C & 8348 \\
 \hdashline
 Chicago & 5300 \\ \hline
 Philadelphia & 1480 \\ \hline
 Washington & 787 \\ \hline
 Baltimore & 115 \\ \hline
 New York & 11
\end{array}"

Draw and label the x and y axes. 

Draw the bars corresponding to the frequencies.

The Pareto chart shows that assaults had the highest frequency and homicides the lowest.

"Total=8348+5300+1480+787+115+11=16041"

"{8348 \over 16041}\cdot100\%=52.04\%"

"{5300 \over 16041}\cdot100\%=33.04\%""52.04\%+33.04\%=85.08\%"

"{1480\over 16041}\cdot100\%=9.23\%""85.08\%+9.23\%=94.31\%"

"{787\over 16041}\cdot100\%=4.91\%""94.31\%+4.91\%=99.21\%"

"{115\over 16041}\cdot100\%=0.72\%""99.21\%+0.72\%=99.93\%"


"{11\over 16041}\cdot100\%=0.07\%""99.93\%+0.07\%=100\%"


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 City & Number & Cum. Relative \ Frequencies (\%)\\ \hline
 D.C & 8348 & 52.04 \\
 \hdashline
 Chicago & 5300 & 85.08 \\ \hline
 Philadelphia & 1480 & 94.31 \\ \hline
 Washington & 787 & 99.21 \\ \hline
 Baltimore & 115 & 99.93 \\ \hline
New \ York & 11 & 100
\\ \hline
 Total= & 16041 & 
\end{array}"

Therefore, the following Pareto Chart is obtained based on the table above: